A ferris wheel has a radius of 7.62m and spins at a constant speed of 0.67m/s. If you were riding in

one of the baskets of the ferris wheel:


How long would it take you to complete one ride of 12 revolutions?


1
Expert's answer
2021-10-26T09:39:40-0400

Let's first find the distance traveled by the rider during one revolution (it equals to the circumference of the wheel):


"C=\pi d=2\pi r=2\pi\times7.62\ m=47.87\ m."

Then, we can find the distance traveled by the rider during one ride of 12 revolutions:


"s=12C=12\times47.87\ m=574.5\ m."

Finally, we can find the time that the rider takes to complete one ride of 12 revolutions:


"t=\dfrac{s}{v}=\dfrac{574.5\ m}{0.67\ \dfrac{m}{s}}=857\ s\times\dfrac{1\ min}{60\ s}=14.3\ min."